Find the current yield of PVR that reported a dividend of $1.88 and closing price of...

  1. Find the current yield of PVR that reported a dividend of $1.88 and closing price of $21.82(rounded to nearest tenth of a percent) ?
  2. A company has 120,000 shares of cumulative preferred stock that pays dividends at $0.25 per share and 200,000 shares of common stock. This year $500,000 is to be distributed. The preferred stockholders are also due to receive dividends in arrears for one year. How much will go to the preferred stockholders for this year's dividend?
  3. Using Table 15-4 find the beginning-of-the-year NAV for the AACNX mutual fund (round to nearest cent).
  4. Find the sales charge percent for one share of the ASCPX mutual fund stock if the stock was offered at $4.94. Use Table 15-4. Round to the nearest hundredth of a percent.
  5. Find the return on investment for 12,500 shares of a mutual fund purchased with an offer price of $16.23 if the shares were sold with a NAV of $16.08. The shares paid a dividend of $0.11 per share during ownership (round to hundredth of a percent).
Answer #1

Answer 1.

current yield = (dividend)/ closing price = [1.88 / 21.82 ] *100 = 8.6%

Answer 2.

Preference dividend = [ 120000 *0.25 ] *2 years = $60000

Answer 5.

Total sale proceeds = 12500 * ($16.08) = $201000

Dividend = 12500 * ($0.11) = $1,375

Total cost purchase = 12500*16.23 = $202875

Gain on investment = ($201000 + $1,375) - $202875 = - 500

ROI = -500/ 202875 * 100 = - 0.25%

Note: for question 3 and 4 tables are required. Please post one question as per policy.